OpenStudy (anonymous):

i must find lim of a_n for x_n = 1/n( 1/ln(2) + ... + 1/ln(n) ) using stoltz cesaro theorem

OpenStudy (anonymous):

xn>int 1/xlnx]1 to inf=ln(lnx)]1 to inf=inf so limxn=inf

OpenStudy (anonymous):

(1/n)*( 1/ln(2) + ... + 1/ln(n) ) is the equation sorry

OpenStudy (anonymous):

I must use stoltz cesaro \[\frac{a_{n+1}-a ^{n} }{ b_{n+1}-b ^{n} }\] for \[x _{n}=\frac{ 1 }{ n}*(\frac{ 1 }{ \ln_{2} }+ ... +\frac{ 1 }{ \ln_{n} })\] and i don t know how to find \[a_{n}\] and \[b _{n}\]

OpenStudy (anonymous):

Please help

OpenStudy (anonymous):

try b_n = n, and a_n = 1/ln2 + ... + 1/ln(n)

OpenStudy (anonymous):

b_n is n or 1/n :-S

OpenStudy (anonymous):

with n get lim = 0 which is good but can i take only n and not 1/n as b_n? :D

OpenStudy (anonymous):

I compare with integral which it approach to infinity then I conclude the series is also infinity or it is diverged.

OpenStudy (anonymous):

i have a hint and says that the result is 0 but if i consider b_n 1/n the result is infinity so i don't know if the hint is good or not or whick one is correct

OpenStudy (anonymous):

ugh you want x_n in the form of a_n/b_n... so tell me, what would a_n be if b_n = 1/n?

OpenStudy (anonymous):

you' re right now i saw that if u write them your way is the same thing, is correct, i understood and the result is 0 as it should be thank you very muc :D i apreciate it